# Recidivism and Peer Influence with LLM Text Embeddings in Low Security Correctional Facilities

### v1

@@ Abstract
We find AI embeddings obtained using a pre-trained transformer-based Large Language Model (LLM) of 80,000-120,000 written affirmations and correction exchanges among residents in low-security correctional facilities to be highly predictive of recidivism. The prediction accuracy is 30% higher with embedding vectors than with only pre-entry covariates. However, since the text embedding vectors are high-dimensional, we perform Zero-Shot classification of these texts to a low-dimensional vector of user-defined classes to aid interpretation while retaining the predictive power. To shed light on the social dynamics inside the correctional facilities, we estimate peer effects in these LLM-generated numerical representations of language with a multivariate peer effect model, adjusting for network endogeneity. We develop new methodology and theory for peer effect estimation that accommodate sparse networks, multivariate latent variables, and correlated multivariate outcomes. With these new methods, we find significant peer effects in language usage for interaction and feedback.
@@ 1 Introduction
Policymakers are particularly interested in predicting and assessing the risk of recidivism (repeat offending) to provide additional remedial measures for prevention, such as psychological interventions (Beaudry et al., 2021). The reconviction rates worldwide are estimated to be high, between 18%-55% across 33 countries (Yukhnenko et al., 2023). Given a worldwide prison population estimate of around 11 million people, the human costs to society of recidivism, therefore, are substantial. In terms of economic costs, a study in 2016 estimated the total costs of incarceration in the US at over one trillion dollars or 6% of GDP, assigning monetary value to different types of costs to individuals (Pettus-Davis et al., 2016).
Psychosocial interventions, including cognitive behavioral therapy (CBT), therapeutic communities (TCs), halfway homes, and psychoeducation, are widely employed in various correctional settings to reduce the risk of recidivism (Beaudry et al., 2021; Lee, 2023; Doleac, 2023). In this paper, we focus on TCs, which are low-security correctional facilities for criminal behavior and substance use disorder (De Leon, 2000). TCs are mutual aid based programs in which residents work together, interact in group therapy as well as informal settings, and are expected to give feedback to peers in the form of affirmations for prosocial behavior and corrections for behavior that contravenes TC norms (De Leon, 2000; Perfas, 2014). Affirmations and corrections can either be informal comments or can be written down on “slips”, brief forms that typically include the sender, the receiver, and the date and content of the affirmation/correction. Notably, Randomized controlled trials have shown TCs to reduce criminal activity and drug use (Sacks et al., 2012; Bahr et al., 2012).
However, assessing the risk of recidivism is challenging (Dressel and Farid, 2018; Lin et al., 2020; Campbell et al., 2019; Warren et al., 2020). Risk is often evaluated using individual-level characteristics in a supervised machine learning (ML) algorithm. Dressel and Farid (2018) shows that the widely used COMPAS algorithm, even with high-dimensional features (137 variables) and a proprietary ML algorithm, can achieve only similar levels of prediction accuracy as non-expert human prediction without any criminal justice knowledge. Beyond prediction, a related strand of research explores the factors that reduce the risk of recidivism, including higher employment opportunities Schnepel (2018); Galbiati et al. (2021), welfare programs Yang (2017); Tuttle (2019), and neighborhood institutions Barrios-Fernández and García-Hombrados (2025).
We find a similar situation in terms of predicting recidivism in our data in this paper as well. Our data comes from 3 TCs in a midwestern state in the United States. The TCs maintained administrative records of each resident’s precise entry and exit dates, pre-entry covariates, time-stamp, sender ID, receiver ID, and the actual text of all the corrections and affirmations that residents exchanged during their time in TC. The midwestern state’s department of rehabilitation and corrections provided the data on 3-year recidivism. Using all pre-entry covariates in our setup, we can only predict 3-year recidivism up to an out-of-sample AUC value slightly above 0.5, i.e., marginally better than random guessing. However, TCs are unique in their structure of peer affirmations and corrections, along with the practice of recording them. This unique feature allows for analyzing interpersonal language exchanges within the program as predictors of critical downstream outcomes, particularly criminal recidivism.
Although the language exchange of residents holds necessary signals for downstream outcomes, we cannot directly use this unstructured data corpus for the prediction exercise. As a first step, we use transformer-based Large Language Models to represent this language as numerical embedding vectors that can be used as input for prediction. We find that we can significantly improve recidivism prediction with an (out-of-sample) AUC of 0.70 using the transformer-based Large Language Models (LLM) embeddings (Vaswani et al., 2017; Devlin et al., 2019) of the repository of 80,000 to 120,000 written affirmations and corrections that the residents exchange as part of the mutual aid-based therapy in TCs. This is a $\approx$ 30% improvement in predicting recidivism over using pre-entry characteristics of residents and comparable to state-of-the-art commercial softwares such as COMPASS that uses orders of hundreds of individual characteristics. This supports recent work in TC that has focused on community building, peer interactions, and peer encouragement as possible mechanisms behind the success of TCs (Campbell et al., 2019; Nath et al., 2025).
After establishing the importance of interactions on recidivism, we explore the role of peers in shaping individuals’ engagement and behavior profiles while they are in the TCs. We obtain estimates of peer effects on their language embeddings to do so. But embeddings are high-dimensional and not interpretable directly. We perform Zero-Shot classification (Yin et al., 2019) to classify the embeddings into interpretable user-defined classes. Here, the classifier generates probabilities for each class supplied by the user that add up to 1. However, there are several challenges to estimating peer effects in this context, including multivariate outcomes, sparse networks, and endogeneity in network formation due to multivariate latent homophily, which the existing methods for peer effect estimation do not accommodate. To overcome these challenges, we develop new econometric methods based on instrumental variables (IV) that accommodate correlated multivariate outcomes, multidimensional latent variables, do not require dense networks, and use flexible nonparametric modeling of the effects of latent homophily. We show our estimator is $\sqrt{N}$ consistent and asymptotically normal.
We contribute to the growing literature on text analysis to address various economic questions. The works of Kalamara et al. (2022); Lippmann (2022); Jelveh et al. (2024); Brehm and Saavedra (2025) use some of the earlier approaches to text analysis, such as sentiment analysis, word frequency counts, bag of words, topic modeling, and the dictionary methods. More recently, Gennaro and Ash (2022); Ash et al. (2022) use the early models of word embedding, such as Word2Vec (Mikolov et al., 2013), as methods for text analysis.
Unlike these papers, we use the embeddings from transformer-based foundation AI models that use the attention mechanism in Vaswani et al. (2017). The excellent expressive ability of the embeddings of transformer-based models, such as GPT, is evident by their exceptional generative and predictive abilities, in the current generative AI revolution (Rathje et al., 2024). This is due to the attention mechanism and bidirectional processing, whereby other words in the sentence dynamically influence the embeddings of words, defining a context for the word. In this aspect, our approach is related to Bajari et al. (2025), who used embeddings to create hedonic price indices using data from Amazon. Further, unlike the aforementioned papers using text data, we move beyond the representation using embeddings and implement Zero-Shot classification to assign the text to interpretable user-defined classes. The Zero-Shot method turns the classification problem into an inference problem by fine-tuning deep representations from the transformer on a specialized language inference task and provides a proximity (entailment) of a given text to an arbitrary user-defined phrase. This entailment inference does not need any new training data, making it well-suited for applications like ours that do not have a large corpus of training data. Our prediction performance on recidivism using the probabilities generated for the user-defined classes (Zero-Shot) retains the predictive power of the results obtained using the high-dimensional text embeddings.
The linear-in-means peer effect model has been widely employed in econometrics for estimating peer effects (Manski, 1993; Bramoullé et al., 2009; Johnsson and Moon, 2021; Hayes and Levin, 2024; Bramoullé et al., 2020). In a seminal paper, Bramoullé et al. (2009) showed conditions under which the peer effect parameter is identifiable, enabling a Two Stage Least Squares (2SLS) solution when the network is not correlated with error conditional on covariates. However, typically in observational data, the network is endogenous, due to common latent variables explaining both network formation and outcomes. Several authors have proposed solutions to account for this unobserved heterogeneity by modeling the network generation process with a latent variable model and estimating the latent variables from observed network data (Goldsmith-Pinkham and Imbens, 2013; Nath et al., 2025; Chang and Paul, 2024; Johnsson and Moon, 2021).
The methodology developed in this paper applies to more general settings of peer effects, as our methods allow for multivariate latent traits to model the network formation relative to a univariate latent homophily variable in Johnsson and Moon (2021); Goldsmith-Pinkham and Imbens (2013), sparse networks, unlike the dense networks required in Auerbach (2022); Graham (2017); Johnsson and Moon (2021), and for multivariate outcomes. Two other papers that develop methods for peer effects in multivariate outcomes are Zhu et al. (2020); Cohen-Cole et al. (2018), but they do not account for endogeneity in network formation. Therefore, our proposed methodology expands the scope of peer effect estimation. Before implementing the new methods on real data, we assess the performance of the new estimator in numerical simulations. This analysis shows that the error rate in recovering all direct and indirect peer effect parameters declines steadily as we increase the sample size.
One key contribution of this paper’s new methodology is the estimation of network formation using the multivariate latent variables and nonparametric modeling of the effect of latent variables on the outcome, regressors, and instruments. As a precursor to the peer effects estimation, we compare the performance of four network models (tetra logit and joint fixed effect MLE from Graham (2017), RDPG model from Athreya et al. (2017), latent space without covariates in Hoff et al. (2002) and finally latent space with latent multiplicative, additive factors and observed covariates in Ma et al. (2020) in explaining three critical network properties: modularity, standard deviation of row means, and transitivity of the clustering coefficient on our real data application. For both the male and female units, we find that the latent space model allowing for additive, multiplicative latent variables and observed covariates performs best in replicating the above-listed network properties.
With this new methodology, we estimate the multivariate peer effects model controlling for network endogeneity in language exchanges within the TC. We find evidence of network peer effects when aggregated by sender or receiver profiles. However, the magnitude of the direct peer influence is much larger than the indirect spillover across both male and female units. Figure B1 (Appendix) summarizes our pipeline from data to the findings and contributions.
@@ 2 Empirical Setup
This paper uses datasets from two male units and one female unit of Therapeutic Communities (TCs) located in a midwestern state in the United States. These TCs were freestanding, minimum security, community-based correctional facilities. Thus, while they were locked facilities, they were not units in a prison. These TCs drew from a mixed urban/suburban/rural catchment area; each unit had 80 beds. Since data was collected over time, the female unit had 982 residents, and the male units had 1649 residents. The two male units are located on two floors of the same TC, and henceforth they will be collectively referred to as the male unit. We observe the entry and exit dates of these residents. The residents may stay in these units for up to 180 days. They may either successfully graduate from the program at the end of their stay or leave it unsuccessfully. We find, on average, the residents stayed for about 120 days. Table B1 (in the Appendix) details the residents used in the analysis. There is a very high graduation rate from these units, with 87% of residents successfully graduating from the female unit, and the corresponding percentage of graduates for the male unit is 89%.
@@
Given that we observe the time stamps of entry and exit, we observe a substantial variation in these dates and the residents’ length of stay in these units. Figure B2 (a,b) shows the variation in these entry and exit dates, and panel (c) shows substantial differences in the length of the stay. The resident information from TCs is mapped with the administrative dataset maintained by the State’s Department of Rehabilitation and Corrections. This administrative dataset provides the date of reincarceration of TC graduates, tracked up to 3 years. Reincarceration is one of several possible measures of recidivism, but it is the most commonly used in the criminology literature (Bouchard and Wong, 2024) and has been used in previous studies of TC outcomes (Doogan and Warren, 2016; Warren et al., 2007, 2020). Table B1 shows that 22% of the residents had recidivism in the female unit and 25% in the male unit (among all residents, irrespective of graduation/success from TC treatment status). These numbers align with global figures obtained by Yukhnenko et al. (2023) across 33 countries and various types of crimes.
@@
During their stay in TCs, the residents were encouraged to exchange written affirmations and corrections with each other as part of their therapy. We observe the timestamps, text, sender, and receiver IDs for each of these exchanges in these TCs. The digitized records of these texts consist of 123,000 such exchanges between the residents in the female unit and about 83,000 exchanges between the residents in the male unit. Figure B3 (Appendix) shows the distribution of the count of affirmations (push) and corrections (pull) aggregated by sender ID in panels (a,b) and by receiver ID in panels (c,d), respectively. The female unit shows differences in the distribution of affirmations and corrections, with affirmations being less skewed. However, we see an almost overlapping distribution of corrections and affirmations in the male units.
@@
We deploy transformer-based LLMs to create text embeddings separately for the 123,000 and 83,000 written exchanges in the female and male units. Precisely, we use the BERT (bert-base-uncased) language model from Devlin et al. (2019). The implementation is done using the text package in R (Kjell et al., 2023) that allows users to access these LLMs easily from huggingface.co. We use these embeddings to create sender and receiver profile vectors for the residents based on the language they use to interact with their peers. These embeddings represent how the behavior and engagement of individuals vary within the TC. The sender profiles indicate how individuals interact with their peers, for example, being supportive or dismissive of others. The receiver profiles, on the other hand, are related to the behavior of the individuals as perceived by their peers. This is because the messages individuals receive reflect their conduct during their stay in the TC (for example, being called out for disruptive conduct). We use these embeddings to assess two key questions- one related to predictive ability for recidivism and the other related to the mechanism of peer influence. We describe in Section 5 that these word embeddings have substantial predictive ability for recidivism.
However, it is hard to interpret these 768 dimensions meaningfully. Therefore, our second method uses a LLM-based Zero-Shot text classifier Yin et al. (2019); Meshkin et al. (2024); Wang et al. (2023) using the BART model of Lewis et al. (2019). In this method, the user provides a fixed class of words to the LLM along with the text messages, and the method generates predictions for the user-supplied classes for each message. Note that the classification is performed without seeing any labeled training data (and hence “Zero-Shot”), and with only the knowledge of natural language that the LLM has acquired.
Figure 2: Raw messages and their predicted class probabilities using transformer-based Zero-Shot classifier
@@
The classification mapping is explained in Figure 2. The figure provides examples of 5 text messages from TCs and their associated LLM-based Zero-Shot classifier probabilities. This method is implemented using the transforEmotion package in R (Christensen et al., 2024). Similar to the text package, any LLM-based Zero-Shot classifier with a pipeline on Huggingface can be implemented on the local machine without sending any data to an external server. The four user-supplied classes are chosen using the prompts to the latest AI models, where we provide a random sample of 1000 written texts and query to provide four short phrases that best describe these messages for Zero-Shot classification. The LLM-based Zero-Shot classification method has been shown to achieve performance comparable with deep neural networks with access to labeled training data in biomedical applications (Meshkin et al., 2024).
We display the scatter plot for class probabilities by message type. The classification appears to work well as the probabilities for “rule violations” and “disruptive conduct” are much higher for corrections, and we see the opposite for affirmations (Figure 3 for females and Figure B4 for males, respectively). We review the methodology behind transformers, LLMs, and Zero-Shot classification in Section 3.
@@
Our final step is to study the underlying behavioral mechanism for these exchanges. Using the low-dimensional interpretable class probabilities, we set up a multivariate outcome model of peer effects that allows for contextual and endogenous peer effects and endogeneity in the network formation. The network links are generated using the information on entry and exit date stamps, along with the dates and frequency of the message exchanges. We describe our peer effect estimation methodology in more detail in section 3.

### v2

@@ Abstract
Studying peer effects in language is critical because they often reflect behavioral and personality traits that are important determinants of economic outcomes. However, language is unstructured, non-numeric, and high-dimensional. We combine Large Language Model (LLM) embeddings with structural econometric identification to provide a unified framework for identifying peer effects in language. This unified framework is applied to 80,000–120,000 written exchanges among residents of low security correctional facilities. The LLM language profiles predict three-year recidivism 30% more accurately than pre-entry covariates alone, showing that text representations capture meaningful signals. We analyze peer effects on multidimensional language embeddings while addressing network endogeneity. We develop novel instrumental variable estimators for peer effects that accommodate multivariate outcomes, sparse networks, and multidimensional latent variables. Our methods achieve $\sqrt{N}$-consistency and asymptotic normality under realistic sparsity conditions, relaxing the dense-network assumption. Results reveal significant peer effects in residents’ language profiles.
JEL Classification: C31, C36, C55, K42.
Keywords: Peer effects, Language embeddings, Network endogeneity, Recidivism.
@@ 1 Introduction
Peer effects play an important role in many outcomes economists are interested in, including educational attainment, criminal behavior, and workplace productivity (Bayer et al.,, 2009; Mas and Moretti,, 2009; Sacerdote,, 2001). Researchers have identified several mechanisms by which peers influence individuals’ behavior (Boucher et al.,, 2024; Bursztyn et al.,, 2014). These include social learning, where individuals learn from or update beliefs based on peers’ actions (Banerjee,, 1992; Bikhchandani et al.,, 1992), and models of social image and peer pressure, where individuals modify behavior to adapt to societal norms (Bursztyn and Jensen,, 2017). In this paper, we focus on language as an underlying mechanism and propose studying peer effects on language. It is well known that individuals often convey behavioral traits and beliefs through language exchanges (Fouka,, 2020; Grogger,, 2019; Giles and Ogay,, 2013; Niederhoffer and Pennebaker,, 2002). Therefore, peers’ language reveals their social alignment and beliefs, making it central to these mechanisms of peer effects.
There are two major challenges to estimating peer effects in language. First, language interactions, by nature, are non-numerical, unstructured, high-dimensional, and context-dependent, making them difficult to analyze mathematically. Second, similar to peer effect estimation in other settings, the identification of peer effects in language is complex due to multiple issues. These include simultaneity in the outcomes and confounding due to unobserved homophily affecting both the outcome and peer selection (Manski,, 1993; Shalizi and Thomas,, 2011; Johnsson and Moon,, 2021).
The first challenge is to represent language interactions in a structured numerical form. Traditional text analysis methods, such as bag-of-words, dictionary methods, and topic models have limitations in capturing implicit semantic meaning and often require explicit appearance of keywords specified by the researcher in the text. Earlier text embedding models, such as Word2vec (Mikolov et al.,, 2013) attempt to address these limitations by representing text as high-dimensional embedding vectors. However, even these models are limited by their inability to incorporate context in their embeddings. In terms of the second challenge, identification of peer effects can be achieved using instrumental variables, where peers interact in network settings, using the identification results outlined in Bramoullé et al., (2009). However, Bramoullé et al., (2009)’s results assume that the network is uncorrelated with the error term conditional on covariates. Recently, Johnsson and Moon, (2021) have extended the results of Bramoullé et al., (2009), allowing for network endogeneity. However, these results rely on strong assumptions about network density that are unrealistic for real-world network data. Consequently, the identification and estimation of peer effects in language require new econometric methods and theory that accommodate multivariate, correlated outcomes, endogeneity in network formation, and are developed under the assumption of realistic, sparse networks.
In this paper, we provide a unified framework that combines Large Language Model embeddings with novel econometric methodology for the identification and estimation of peer effects in language interactions. We use embeddings from transformer-based foundation AI models that improve upon early text embedding models by incorporating both an attention mechanism (Vaswani et al.,, 2017) and bidirectional processing (Devlin et al.,, 2019) to obtain context-aware vector representations of text. These models can generate different vector embeddings for the same word because other words in the sentence dynamically influence its embeddings, thereby defining a context for the word. Foundation models trained on extensive text corpora have demonstrated strong capabilities for learning language representations. Consequently, these models can be applied to obtain embeddings for new text in specific applications without the need for any new training (Radford et al.,, 2018; Brown et al.,, 2020).
Even though the LLM embedding vectors provide a meaningful numerical representation of language, they are not well-suited for peer effect estimation due to their high dimensionality (700-1000). Such high dimensionality limits the scope of interpretation in peer effect estimation. We address this problem using zero-shot classification, which provides a suitable way to map high-dimensional text embeddings to interpretable user-defined input label categories. Zero-shot classification reframes the classification task as a language inference problem that determines entailment between a text and the user-supplied label categories (Yin et al.,, 2019). Notably, this language inference task does not require labeled training data. On the peer effect identification side, we develop new econometric methods based on instrumental variables (IV) to accommodate correlated multivariate outcomes and multidimensional latent variables. Our methods do not require dense networks and accommodate flexible nonparametric modeling of multidimensional latent homophily vectors. The proposed estimator is shown to be $\sqrt{N}$ consistent and asymptotically normal.
We apply this new unified framework of LLM embeddings and econometric methodology to written exchanges between residents of three low-security correctional facilities in a midwestern state in the United States. These correctional facilities, formally known as therapeutic communities (TC), are for individuals with criminal behavior and substance use disorder (De Leon,, 2000). These programs are based on mutual aid, encouraging residents to be collaborative, practice group therapy, and provide regular feedback to their peers using affirmations and corrections. Corrections are targeted towards individuals for violating TC norms, and affirmations are given to encourage positive and prosocial behavior (De Leon,, 2000; Perfas,, 2014). Affirmations and corrections are written on “slips,” which are brief forms containing fields for the sender, receiver, and date, along with the message. Previous randomized controlled trials have demonstrated that TCs reduce criminal activity and drug use (Sacks et al.,, 2012; Bahr et al.,, 2012).
In these three correctional facilities, we observe the administrative records for each resident’s precise entry and exit dates, pre-entry covariates, and details on the content, date, sender, and recipients of these written exchanges. These residents are then mapped with the recidivism records provided by the midwestern state’s department of rehabilitation and correction. These correctional facilities are unique both in their practice of exchanging written affirmations and corrections and also in their process of recording them. This unique feature enables us to study peer effects in these language interactions and to use information contained in language to predict meaningful downstream outcomes, such as recidivism.
Using only the pre-entry covariates, we can predict 3-year recidivism with an out-of-sample AUC only slightly above 0.5, which is just marginally better than random guessing. However, using transformer-based LLM embeddings of the 80,000 to 120,000 written exchanges, prediction accuracy improves dramatically to an AUC of 0.70, approximately 30% higher than using pre-entry characteristics alone. This performance is comparable to or slightly better than published AUC values from state-of-the-art commercial software, such as COMPAS, which uses hundreds of individual characteristics in a machine learning model, raising concerns about fairness (Dressel and Farid,, 2018). An AUC of 0.70 is also generally considered “good” in the context of predicting criminal recidivism (Laqueur and Copus,, 2024). These results establish that LLM representations of written exchanges capture economically meaningful behavioral variation. This finding further supports recent work suggesting that peer interactions are key mechanisms behind TC effectiveness (Campbell et al.,, 2019; Nath et al.,, 2025).
After establishing that written exchanges in TCs contain meaningful signals for recidivism, we examine whether these peers influence each other’s language profiles. Using zero-shot classification, the unstructured language exchanges are mapped to four user-defined input label categories that represent distinct behavioral dimensions. These chosen input labels are community support, personal growth, rule violations, and disruptive conduct. [^1] These four input labels for zero-shot classification were chosen using principled, iterative, and replicable LLM prompts; the steps are detailed in Appendix A3. Applying the peer effects methodology developed in this paper to the predicted probabilities for language categories reveals significant peer effects. Residents affect each other’s language profiles even after accounting for unobserved homophily using the new methods developed in this paper. Peer effects within the same language dimension are substantially larger than cross-category spillovers across dimensions, a pattern observed consistently across both male and female units and across sender and receiver profiles.
The novel econometric methods developed and employed in this study indicate significant peer effects in language profiles, and predictive analysis demonstrates that language exchanges are associated with 3-year recidivism. Nevertheless, predictive power alone is not sufficient for policy implications. Estimating the causal impact of language profiles on recidivism is necessary to determine whether language exchanges in TC can causally affect three-year recidivism. The primary identification challenge in this causal relationship is the presence of unobserved confounders that may influence both language profiles during the stay in TC and recidivism. We estimate the association between language profiles and recidivism, controlling for covariates that we can observe. These covariates include Level of Service Inventory-Revised (LSI-R), which is an aggregate measure of 54 questions provided to participants spanning criminal history, mental health, education, employment, family, living conditions, and alcohol and drug abuse. However, we acknowledge the limitation of this estimate due to unmeasured confounders and use two approaches (Oster,, 2019; Cinelli and Hazlett,, 2020) to bound the magnitude of omitted-variable bias in the discussion section. We find that estimates of the effect of language in TC on recidivism remain statistically significant at conventional levels, even after accounting for considerable omitted-variable bias relative to the observables.
2 Related Literature and Our Contributions
Text analysis is becoming increasingly important in economics for extracting meaningful insights from unstructured data. The works of Kalamara et al., (2022); Lippmann, (2022); Brehm and Saavedra, (2025); Alsan et al., (2025) use some of the earlier approaches to text analysis, such as sentiment analysis, word frequency counts, bag-of-words, topic modeling, and dictionary methods. Some recent works, such as Gennaro and Ash, (2022); Ash et al., (2025), have improved on these earlier models and instead deployed word embedding methods such as Word2Vec (Mikolov et al.,, 2013), for text analysis. However, as outlined in the introduction, these pre-transformer word embedding methods do not account for context. In contrast, we deploy transformer-based foundation models to estimate text embeddings and then use them for zero-shot classification, leveraging the attention mechanism of Vaswani et al., (2017). In this regard, our approach is related to Bajari et al., (2025), which uses embeddings from transformer models to construct hedonic price indices from Amazon data.
The linear-in-means peer effect model has been widely employed in econometrics for estimating peer effects (Manski,, 1993; Bramoullé et al.,, 2009). Bramoullé et al., (2009) established identification conditions leveraging the network structure and instrumental variables when networks are exogenous conditional on observables. However, networks are typically endogenous due to the phenomenon of latent homophily. These are unobserved characteristics that drive both network formation (peer selection) and the outcomes. Recent work addresses this problem through latent variable models (Goldsmith-Pinkham and Imbens,, 2013; Johnsson and Moon,, 2021), but existing methods require unrealistically dense networks (expected degree scaling as $N$) and rely on univariate latent homophily. Other work accommodates multivariate correlated outcomes (Zhu et al.,, 2020; Cohen-Cole et al.,, 2018) but assumes exogenous network formation. We extend the peer effects methodology to simultaneously accommodate sparse networks (expected degree scaling as $\sqrt{N}$), multidimensional latent homophily, and multivariate outcomes, proving $\sqrt{N}$-consistency under realistic sparsity conditions and using nonparametric sieve methods to control for latent confounding flexibly.
We also contribute to modeling networks using multidimensional additive and multiplicative latent-variable models for peer-effect estimation. While the works of Goldsmith-Pinkham and Imbens, (2013); Johnsson and Moon, (2021) integrated network formation with peer effects estimation, these approaches have used simplistic network models with univariate additive latent factors to account for the unobserved confounders. We advocate the use of Random dot product graph (RDPG) models (Athreya et al.,, 2018), which are multiplicative latent variable models, and the latent space models, which accommodate additive and multiplicative latent effects along with observed covariates (Hoff et al.,, 2002; Ma et al.,, 2020; Li et al.,, 2023). In contrast to the one-dimensional latent-variable model of Graham, (2017) that Johnsson and Moon, (2021) employed, the models we deploy use multidimensional latent variables to model network data flexibly. We build methodological and theoretical tools needed to enable the use of this rich class of additive and multiplicative latent space models in peer effect identification, using estimated latent positions with a nonparametric sieve adjustment.
We empirically compare five network specifications (RDPG, latent space with and without covariates, and tetrad logit and joint fixed effect estimators of Graham, (2017)) in their ability to replicate observed network properties (modularity, degree heterogeneity, transitivity) of our correctional facility network data. We find that the latent space model with additive and multiplicative factors plus covariates (Ma et al.,, 2020) best fits our data.
Our findings have important policy implications for the criminal justice system. Recidivism imposes substantial costs. Reconviction rates range from 18-55% across countries (Yukhnenko et al.,, 2023), with U.S. incarceration costs exceeding one trillion dollars annually (Pettus-Davis et al.,, 2016) estimated in 2016. Recidivism prediction has been a central focus in criminal justice policy, with widely-used risk assessment tools such as COMPAS incorporating over 100 pre-entry features (Dressel and Farid,, 2018). Despite extensive feature engineering, prediction accuracy remains limited (Dressel and Farid,, 2018; Lin et al.,, 2020). Our results demonstrate that language exchanges contain rich behavioral signals in correctional settings where pre-entry covariates have extremely limited signals for recidivism. [^2] While prediction tools inform risk assessment and resource allocation, designing effective interventions requires identifying factors that causally affect recidivism. A separate strand of research has identified several such factors, including employment opportunities, welfare programs, and neighborhood institutions (Barrios-Fernández and García-Hombrados,, 2025; Tuttle,, 2019; Galbiati et al.,, 2021).
@@ 3 Empirical Setup
This paper uses datasets from two male units and one female unit of Therapeutic Communities (TCs). These TCs in a midwestern state in the United States were minimum-security, community-based correctional facilities. Thus, while they were locked facilities, they were not units in a prison. These TCs had residents from a mixed urban/suburban/rural area; each unit had 80 beds. Since data was collected over time, the female unit had 982 unique residents, and the male units had 1649 unique residents. The two male units are located on two floors of the same TC, and henceforth they will be collectively referred to as the male unit. We observe the entry and exit dates of these residents. The residents may stay in these units for up to 180 days. They may either successfully graduate from the program at the end of their stay or leave it unsuccessfully. We find, on average, the residents stayed for about 120 days. Table A2.1 (in the Appendix) details the residents used in the analysis. There is a very high graduation rate from these units, with 87% of residents successfully graduating from the female unit, and the corresponding percentage of graduates for the male unit is 89%.
@@
Given that we observe the time stamps of entry and exit, we observe a substantial variation in these dates and the residents’ length of stay in these units. Figure A2.1 (a,b) shows the variation in these entry and exit dates, and panel (c) shows substantial differences in the length of the stay. The resident information from TCs is mapped with the administrative dataset maintained by the State’s Department of Rehabilitation and Corrections. This administrative dataset provides the date of reincarceration of TC residents, tracked up to 3 years. Reincarceration is one of several possible measures of recidivism, but it is the most commonly used in the criminology literature (Bouchard and Wong,, 2024) and has been used in previous studies of TC outcomes (Doogan and Warren,, 2016; Warren et al.,, 2007, 2020). Table A2.1 shows that 22% of the residents had recidivism in the female unit and 25% in the male unit (among all residents, irrespective of graduation/success from TC treatment status). Interestingly, these numbers align with global figures obtained by Yukhnenko et al., (2023) across 33 countries and various types of crimes.
@@
During their stay in TCs, the residents were encouraged to exchange written affirmations and corrections with each other as part of their therapy. We observe the timestamps, text, sender, and receiver IDs for each of these exchanges in these TCs. The digitized records of these texts consist of 123,000 such exchanges between the residents in the female unit and about 83,000 exchanges between the residents in the male unit. Figure A2.2 (Appendix) shows the distribution of the count of affirmations (push) and corrections (pull) aggregated by sender ID in panels (a,b) and by receiver ID in panels (c,d), respectively. The female unit shows differences in the distribution of affirmations and corrections, with affirmations being less skewed. However, we see an almost overlapping distribution of corrections and affirmations in the male units.
@@
We deploy transformer-based LLMs to create text embeddings separately for the 123,000 and 83,000 written exchanges in the female and male units. Precisely, we use the BERT (bert-base-uncased) language model from Devlin et al., (2019).  [^3] The implementation is done using the text package in R (Kjell et al.,, 2023) that allows users to access these LLMs easily from huggingface.co. We use these embeddings to create sender and receiver profile vectors for the residents based on the language they use to interact with their peers. These embeddings represent how the behavior and engagement of individuals vary within the TC. The sender profiles indicate how individuals interact with their peers, for example, being supportive or dismissive of others. The receiver profiles, on the other hand, are related to the behavior of the individuals as perceived by their peers. This is because the messages individuals receive reflect their conduct during their stay in the TC (for example, being called out for disruptive conduct). We use these embeddings to assess two key questions- one related to predictive ability for recidivism and the other related to the mechanism of peer influence. We describe in Section 6 that these word embeddings have substantial predictive ability for recidivism.
However, it is hard to interpret these 768-dimensional embeddings meaningfully. Therefore, our second method uses a LLM-based Zero-Shot approach (Yin et al.,, 2019; Meshkin et al.,, 2024; Wang et al.,, 2023) using the BART model of Lewis et al., (2019). In this method, the user provides a fixed class of words to the LLM along with the text messages, and the method generates predictions for the user-supplied classes for each message. Note that the classification is performed without seeing any labeled training data (and hence “Zero-Shot”), and with only the knowledge of natural language that the LLM has acquired.
Figure 2: Raw messages and their predicted class probabilities using Zero-Shot classifier
@@
The classification mapping is explained in Figure 2. The figure provides examples of 5 text messages from TCs and their associated LLM-based Zero-Shot classifier probabilities. [^4] This method is implemented using the transforEmotion package in R (Christensen et al.,, 2024). Similar to the text package, any LLM-based Zero-Shot classifier with a pipeline on Huggingface can be implemented on the local machine without sending any data to an external server. The four user-supplied classes are selected using LLM prompts to the latest AI models in a principled, iterative, and reproducible manner. See section A3 in the Appendix for further details on the input label generation. In summary, for choosing the input labels, we generate 100 random samples of text messages stratified by message type. These random samples are then supplied to the Claude API (claude-sonnet-4) with a fixed query for 100 iterations. Finally, the input labels are chosen based on the clusters that well represent the LLM-generated labels in the 100 iterations. Notably, the LLM-based Zero-Shot classification method has been shown to achieve performance comparable to that of deep neural networks with access to labeled training data in biomedical applications (Meshkin et al.,, 2024).
We display the scatter plot for class probabilities by message type. The classification appears to work well as the probabilities for “rule violations” and “disruptive conduct” are much higher for corrections, and we see the opposite for affirmations (Figure 3 for females and Figure A2.3 for males, respectively). We review the methodology behind transformers, LLMs, and Zero-Shot classification in Section 4.
@@
Our final step is to study the underlying behavioral mechanism for these exchanges. Using the low-dimensional interpretable class probabilities, we set up a multivariate outcome model of peer effects that allows for contextual and endogenous peer effects and endogeneity in the network formation. The network links are generated using information on entry and exit date stamps, along with the dates, sender ID, receiver ID, and message exchange frequency. We describe our peer effect estimation methodology in more detail in section 4.